OpenStudy (anonymous):

Show the identities you use to evaluate the following. No calculators allowed! b. cot(π/2 – x) if tan x = -5

OpenStudy (ajprincess):

\[\cot(\frac{ \pi }{ 2 }-x)=\frac{ 1 }{ \tan(\frac{ \pi }{ 2 }-x)}\] \[\tan(\frac{ \pi }{ 2 }-x)=\cot x\] \[\frac{ 1 }{ \cot x }=\tan x\] Does that help? @sfb